OpenStudy (anonymous):

Please Help: Write an equation of the line that passes through (6, -10) and is perpendicular to the line that passes through (4, -6) and (3,-4)?

OpenStudy (anonymous):

The product of the slopes of two perpendicular lines is -1. So, first find the slope of the line that passes through the two points given using the formula slope = (y1-y2)/(x1-x2) = (-6+4)/(4-3) = -2 slope of the line perpendicular to this one will b m x -2 = -1, -2m = -1, m = 1/2 So, the slope of the other line is: (y+10)/(x-6) and that = 1/2 so we have 2y + 20 = x-6 2y = x - 26 y = x/2 -13

OpenStudy (anonymous):

thank you so much.